What is the vertex of the function g(x)=(x+5)2? I need a quick answer please
Answer by stanbon(75887)

What is the vertex of the function g(x)=(x+5)2?
Form:
y-0 = (x+5)^2
----
Vertex: (-5,0)
===================
Cheers,
Stan H.
=====================